an amount of money is invested @ 4% annual simple interest, and twice the amount is invested at 5%. The total annual interest is $112. How much is invested at each rate?
Answer by checkley77(12569)

.05*2x+.04x=112
.10x+.04x=112
.14x=112
x=112/.14
x=800 is the amount invested @ 4%
2*800=1600 invested @ 5%.
Proof:
.05*1600+.04*800=112
80+32=112
112=112